OpenStudy (anonymous):

Write the following numbers in the form re^(iθ) and in the form a + ib: (1+i)^20

OpenStudy (freckles):

let x=1+i x^2=1+2i+i^2=2i and \[x^{20}=x^{2 \cdot 10} =(x^2)^{10}=(2i)^{10}\]

OpenStudy (freckles):

can you simplify it from there ?

OpenStudy (anonymous):

Yeah, the part I don't understand us writing it in re^(iθ) form

OpenStudy (anonymous):

re^(i*theta)

OpenStudy (freckles):

so you got what after computing (2i)^10?

OpenStudy (freckles):

because we can use that to write in re^{i theta) form

OpenStudy (anonymous):

-1024

OpenStudy (freckles):

cos equals -1 for what angle?

OpenStudy (anonymous):

cos(pi)=-1

OpenStudy (freckles):

-1024(1+0i)=1024(-1+0i)

OpenStudy (freckles):

yes

OpenStudy (freckles):

so \[-1024(1+0i)=1024(-1+0i)=1024(\cos(\pi)+i \sin(\pi))\]

OpenStudy (freckles):

and guess what that equals :)

OpenStudy (anonymous):

-1024

OpenStudy (freckles):

well in re^(i theta) form

OpenStudy (freckles):

remember in earlier threads we said \[e^{i \theta}=\cos(\theta)+i \sin(\theta)\]

OpenStudy (anonymous):

1024e^(pi*i)

OpenStudy (freckles):

bring

OpenStudy (freckles):

bingo

OpenStudy (freckles):

I think I'm getting tired i keep typing wrong things

OpenStudy (freckles):

bring lol

OpenStudy (anonymous):

Thank you for all your help tonight! Its okay lol me too.

OpenStudy (freckles):

did this all make sense

OpenStudy (freckles):

i sorta used the method we used on last thread to find that binomial to the 20th power

OpenStudy (freckles):

i like it :)

OpenStudy (freckles):

when it can be used anyways

OpenStudy (anonymous):

Yeah it did make sense! it will help a lot for the other problems like it :)

OpenStudy (freckles):

so the longer way \[(1+i)^{20} =(\sqrt{2}(\frac{\sqrt{2}}{2}+\frac{\sqrt{2}}{2}i))^{20}=(\sqrt{2})^{20}(\cos(\frac{\pi}{4})+\sin(\frac{\pi}{4})i)^{20} \\ = 2^{10}(\cos(20 \frac{\pi}{4})+i \sin(20 \frac{\pi}{4})) =2^{10}(\cos(5\pi)+i \sin(5 \pi))=2^{10}(-1+0)\]

OpenStudy (anonymous):

ick

OpenStudy (anonymous):

i thought the first way was rather snappy

OpenStudy (freckles):

yes that one was more the bread pudding way and this other way was like icky odd yogurt

OpenStudy (anonymous):

then again i guess \[\huge 1+i=\sqrt{2}e^{\frac{\pi}{4}i}\] will work nicely too

OpenStudy (anonymous):

no not really since \(\frac{\pi}{4}\times 20=5\pi\) it is not so bad

OpenStudy (anonymous):

i have to admit i would never have thought to compute \((1+i)^2\) first